OpenStudy (anonymous):

if f(x)=3x^2-8x^-2 then lim as h approaches 0 f(2+h)-f(2)/h=

OpenStudy (anonymous):

Thats the derivative of f at x = 2 f'(x) = 6x - 8 f'(2) = 4 It is 4.

OpenStudy (ness9630):

Solve using the quotient rule right?

OpenStudy (anonymous):

It is not 4. It's a multiple choice question and my choices are: 10, 14, 20, -14, and -20.

OpenStudy (campbell_st):

you are being asked to use 1st princpals to find the derivative so this is tedious.... but do-able.. so substiute 2 + h into f(x) or \[f(2 + h) = 3(2 + h)^2 - \frac{1}{(2 + h)^2} \] and f(2) = 12 - 1/4 = 47/4

OpenStudy (campbell_st):

so you are looking at \[\lim_{h \rightarrow 0} \frac{(2 + h)^2 - \frac{1}{(2 + h)^2} - \frac{47}{4}}{h}\] all you need to do now is simplify

OpenStudy (campbell_st):

good luck

OpenStudy (anonymous):

Thank you!

OpenStudy (campbell_st):

but I'm use the cheats method find \[f'(x) = 6x + \frac{16}{x^3}\] then find f'(2) for the correct answer

OpenStudy (campbell_st):

oops 1'd not 1'm